OpenStudy (anonymous):

Is (1, 5) a solution to the equation y=x-2?

OpenStudy (anonymous):

plug it in. (5) = (1) - 2 are the sides equaivalent?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that means that (1,5) is niot a solution for y=x-2.

OpenStudy (anonymous):

Saying that (1,5) is not a point on y=x-2.

OpenStudy (anonymous):

ok so for if i had (-2,-6) asking if it was a solution to y=3x? (-6)=(-2)-3x? that wouldn't be a solution etiher?